the standard enthalpy of combustion of ethane, carbon, and hydrogen are -1559.7 kJ/mol, -393kJ/mol and -286kJ/mol respectively.
Use Hess'ss law to find the enthalpy of formation of ethane
1
Expert's answer
2019-12-11T07:54:29-0500

ΔHform(C2H6)=ΔHcomb(C2H6)-(2ΔHcomb(C)+3ΔHcomb(H2))=-1559.7-(2(-393)+3(-286))=84,3 kJ/mol